I need to graph the following equation and indentify the y-intercept
y+x=-7
Thank you

Answer by rfer(16322)

y int is when x=0
y+0=-7
y=-7
y int=(0,-7)
-----------
second point x int
0+x=-7
x=-7
x int=(-7,0)